# TriNeRFLet: A Wavelet Based Multiscale Triplane NeRF Representation

### v1

@@ 3 TriNeRFLet
4 TriNeRFLet Super-Resolution
Given our TriNeRFLet scheme, we present a novel NeRF super-resolution approach that relies on the TriNeRFLet special multiscale structure. This method employs a 2D pre-trained stable diffusion (SD) upscaler [33] for super-resolution ”guidance”. Our approach does not require any kind of low-high resolution 3D scene pairs for supervision as done in other works [4]. Utilizing the robustness of the SD upscaler, our approach has the potential to handle different types of scenes and a range of resolutions.
Given a set of low-resolution (LR) multiview images of a given scene, the goal is to generate high-resolution novel views for this scene. We fit a TriNeRFLet to the LR images and aim at estimating a high-resolution (HR) TriNeRFLet. Denote the size of the LR planes by $N_{LR}$ and the number of wavelet levels by $L_{LR}$. For the HR planes, denote the size by $N$ and the number of levels by $L$. Due to the multiscale property of the wavelet representation, in the HR TriNeRFLet we just need to learn the high frequencies of the wavelet representation, i.e., the first $L_{LR}$ levels of the HR wavelet planes are shared with the LR TriNeRFLet.
$N_{LL}$
$L$
$N_{base}$
$N_{final}$
$C$
$\gamma$
$W$
$D_{dens}$
$D_{col}$
train steps
Small
$64$
$4$
$512$
$1024$
$16$
$0.2$
$64$
$1$
$2$
$6k$
Base Light
$64$
$5$
$512$
$2048$
$32$
$0.4$
$64$
$1$
$2$
$10k$
Base
$64$
$5$
$512$
$2048$
$32$
$0.4$
$64$
$1$
$2$
$43k$
Large
$64$
$5$
$512$
$2048$
$48$
$0.6$
$128$
$1$
$2$
$83$
Table 1: The parameters used in the different TriNeRFLet versions
@@ 4 TriNeRFLet Super-Resolution
Figure 5 illustrates our TriNeRFLet super-resolution strategy. First, we start by fitting a LR TriNeRFLet to the provided low-resolution multiview images. After training for $s_{LR}$ steps, the super-resolution process starts, and a key component of it is the stable diffusion upscaler refinement ($\text{SD}_{\text{refine}}$) step presented in Figure 6. Given a low resolution ground truth image $\bm{x}^{gt}_{LR}$, its high resolution version $\bm{x}^{est}_{HR}$ is rendered using all wavelet levels $L$ using the TriNeRFLet. Initially, the result will not be of high-resolution. Thus, to improve it we use a diffusion step with time $t$ randomly selected from the range $[T_{min}\;,\;T_{max}]$. Then, similar to [14], a noise with variance that depends on $t$ is added to $\bm{x}^{est}_{HR}$. Then we plug the noisy version of $\bm{x}^{est}_{HR}$ into the SD upscaler to perform the diffusion process from step $t$ until the end, while being conditioned on the $LR$ image $\bm{x}^{gt}_{LR}$. This results is the enhanced version of $\bm{x}^{est}_{HR}$, denoted as $\bm{x}^{enhanced}_{HR}$. We denote this enhancement process by $\text{SD}_{\text{refine}}$.
@@
We use the SD upscaler with a small modification compared to the original work [33]. In their work, the denoising step is $\bm{x}_{t-1}=\epsilon(\bm{x}_{t},\bm{z},\varnothing)+\alpha\left(\epsilon(\bm{x}_{t},\bm{z},\bm{y})-\epsilon(\bm{x}_{t},\bm{z},\varnothing)\right)$, where $\epsilon$, $\alpha$, $\bm{z}$ and $\bm{y}$ are the U-net denoiser, a guidance factor, the low-resolution image and the guidance text respectively. Instead of doing the guidance in the text direction, we do it in the low-resolution image direction and without text, i.e., $\bm{x}_{t-1}=\epsilon(\bm{x}_{t},\varnothing,\varnothing)+\alpha\left(\epsilon(\bm{x}_{t},\bm{z},\varnothing)-\epsilon(\bm{x}_{t},\bm{z},\varnothing)\right)$.
@@
Having the refined HR images, we want to use them to update the TriNeRFLet. We add these images to the set denoted HR-set. This set is initialized to be empty and gets updated sequentially. It is refreshed (i.e. emptied) every $s_{refresh}$ steps as illustrated in Algorithm 1. The images in HR-set are used as ”ground-truth” images for training the HR TriNeRFLet to update all its $L$ levels.
@@
Method
Mic
Chair
Ship
Materials
Lego
Drums
Ficus
Hotdog
Avg.
Train Time $\downarrow$
Render FPS $\uparrow$
Vanilla NeRF
$32.91$
$33.00$
$28.65$
$29.62$
$32.54$
$25.01$
$30.13$
$36.18$
$31.005$
$\geq 1\text{ day}$
$\leq 0.2$
INGP
$36.22$
$35.00$
$31.10$
$29.18$
$36.39$
$26.02$
$33.51$
$37.40$
$33.176$
$\sim 5\text{ mins}$
$0.6-1.2$
3D Gaussian Splatting
$35.36$
$\underline{35.83}$
$30.80$
$\underline{30.00}$
$35.78$
$\underline{26.15}$
34.87
$\underline{37.72}$
$33.32$
$\sim 10\text{ mins}$
$135$
Triplane
$33.85$
$32.83$
$29.58$
$28.15$
$34.70$
$24.86$
$30.35$
$35.80$
$31.26$
$60$-$180$ mins
$1-4$
Ours: TriNeRFLet Small
$35.18$
$33.76$
$30.33$
$29.14$
$35.32$
$25.66$
$33.34$
$36.24$
$32.37$
$12$-$15$ mins
$1-4$
Ours: TriNeRFLet Base Light
$35.77$
$35.00$
$31.10$
$29.35$
$36.44$
$25.98$
$33.96$
$36.93$
$33.07$
$60$-$90$ mins
$1-4$
Ours: TriNeRFLet Base
$\underline{36.72}$
$35.54$
$\underline{31.77}$
$29.72$
$\underline{36.66}$
$26$
$33.7$
$37.31$
$\underline{33.43}$
$5$-$8$ hours
$1-4$
Ours: TriNeRFLet Large
37.22 36 32.7 30.34 37.32 26.2
$\underline{34.47}$
37.89 34.017
$\sim 1$ day
$1-4$
Table 2: Blender Dataset PSNR ($\uparrow$) results with train time and render FPS for each method. Bold is best, underline is second.
The TriNeRFLet multiscale wavelet structure plays a crucial role in our SR scheme. First, its multiscale structure establishes a profound connection between high-resolution and low-resolution renderings by forcing them to share information (using the same first $L_{LR}$ wavelet levels). Second, wavelet regularization, especially in the layers that only the HR TriNeRFLet uses, constrains the optimization process and forces the HR TriNeRFLet to learn only useful wavelet coefficients that contribute to HR details. Thus, this unique combination of multiscale wavelet and diffusion models allows the method to learn the high-resolution details it exactly needs without redundancy and artifacts.
The diffusion SR model that we use only supports SR from $128$ to $512$. However, we want to be able to work with other resolutions, such as $100$ to $400$ or $200$ to $800$. For lower resolutions, e.g., going from $100$ to $400$, we pad $\bm{x}^{gt}_{LR}$ and $\bm{x}^{est}_{HR}$ with zeros in order to achieve the desired resolution before plugging them into the SD upscaler. When it finishes, the result is un-padded to the original resolution. For higher resolutions, e.g., going from $200$ to $800$, we randomly crop $128$ and $512$ resolution crops from $\bm{x}^{gt}_{LR}$ and $\bm{x}^{est}_{HR}$, respectively, that correspond to the same relative location. We plug the images into the upscaler, and the result is actually an enhanced version of the crop from $\bm{x}^{est}_{HR}$. Then we only fit the high-resolution render to this crop instead of the whole image. When the HR-set is refreshed, a different random crop will be chosen, thus covering more areas as the process continues.
5 Experiments
We turn to present the results of our method. More visual results appear in the project page (preferred to open in Chrome browser). An ablation study appears in Appendix A.
5.1 3D Scene Reconstruction
For evaluating TriNeRFLet in the 3D recovery task, we define four versions of TriNeRFLet - small, base light, base and large. The parameters of TriNeRFLet are wavelet LL component resolution - $N_{LL}$, base resolution - $N_{base}$, wavelet layers - $L$, final resolution - $N_{final}$, number of channels - $Ch$, wavelet regularization $\gamma$, MLP dim (neurons) - $W$, MLP hidden layers - $D_{density}$ and $D_{color}$ and training steps. These parameters are provided in Table 1 for each model. In the small, base light and base versions, TriNeRFLet uses the exact same MLPs as in INGP. TriNerfLet is trained using the Adam optimizer with a learning rate of $0.01$ and an exponential decay scheduler. The wavelet low frequencies in LL (of size $N_{LL}\times N_{LL}$) are randomly initialized, and the other frequencies in LH, HL and HH at all levels are initialized with zeros. The other training details are very similar to INGP.
Since the wavelet reconstruction component affects the runtime of TriNeRFLet, we start the training with lower resolutions of the planes, training coarse to fine. The initial plane resolution is set to $N_{base}$ and it is increased throughout the training till it reaches the size $N_{final}$ that has $L$ wavelet levels. As described above, due to the wavelet multiscale structure, when we move to the next resolution in training we simply add another level (of higher frequencies) to the wavelet representation and add it to the training, where the lower levels are kept as is (but continue to train).
TriNeRFLet implementation is built upon [39], which is a Pytorch-based implementation of INGP.
Triplane TriNeRFLet Ground-Truth
To evaluate the performance of our method, we use the Blender dataset from [27], which contains $8$ synthetic 3D scenes with complex geometry and lighting. We compare TriNeRFLet with the vanilla NeRF [27], regular Triplane (with base config), INGP [28] and 3D Gaussian Splatting [20]. Results are reported in Table 2, and qualitative examples are presented in Figure 7. We ran all TriNeRFLet versions on a single A6000 GPU. The reported results indeed demonstrate the improvement in performance that the multiscale wavelet brings, as it outperforms regular Triplane by a significant margin, and improves over the current SOTA.
Regarding rendering time, the TriNeRFLet (all versions) frames per second (FPS) are approximately $1-4$, while INGP and 3D Gaussian Splatting are approximately $0.6-1.2$ and $135$ respectively. These FPS are tested on a single A6000 GPU. Note that TriNeRFLet rendering FPS is compatible and even slightly better than INGP, which is known to be fast within the NeRF approaches. 3D Gaussian Splatting, which is not a NeRF method, is faster in its rendering time since it uses a different internal structure than NeRF. When compared only to NeRF-based solutions, our method is very fast with competitive reconstruction performance. Note that although our rendering time is faster, the training time of TriNeRFLet is higher than INGP. This is due to wavelet reconstruction, which needs to be done at each training step, while only once during rendering time.
5.2 Blender Scene Super-Resolution
We turn to present our results for TriNeRFLet super-resolution. The objective here is to generate novel views with ×$\times 4$ resolution as input images. To this end, we tested two different settings on the Blender dataset [27], $100$ to $400$ and $200$ to $800$. For the first one, we use TriNeRFLet with $N_{LR}=256$ and $N=1024$, while in the second setting $N_{LR}=512$ and $N=2048$. In both settings $Ch=16$. We use [12] as a base for implementing TriNeRFLet super-resolution, as it is more suitable for this task.
We compare our method with NeRF*, NeRF-Bi, NeRF-Liif, NeRF-Swin, NeRF-SR (SS) (these methods do not require 3D supervision and are described in [40]) and NVSR (results of $200\rightarrow 800$ were provided to us by the authors) [4]. Results are reported in Table 3. Note that NVSR requires 3D supervision (paired LR and HR multiview images) and therefore it is considered only as a reference. When compared only to methods that use 2D supervised methods, we perform the best in most of the cases. Qualitative examples of our approach appear in Figure 8. More super-resolution results appear in the sup. mat, including results on LLFF dataset.
$100\rightarrow 400$ $200\rightarrow 800$
Method
@@ 5.2 Blender Scene Super-Resolution
NeRF*
$25.56$
$0.170$
$27.47$
$0.128$
NeRF-Bi
$24.74$
$0.244$
$26.67$
$0.175$
NeRF-Liif
$25.36$
$0.125$
$27.34$
$0.096$
NeRF-Swin
$24.85$
$0.108$
$26.78$
$0.086$
NeRF-SR (SS)
$\underline{28.07}$
$\underline{0.071}$
$28.46$
$0.076$
TriNeRFLet-SR
28.55 0.061
$\underline{29.49}$
$\underline{0.051}$
NVSR (3D supervised)
***** * 29.53 0.045
Table 3: Blender ×$\times 4$ Super Resolution Results on low resolution input of $100\times 100$ and $200\times 200$. NVSR is trained with multiview (3D) supervision and therefore we consider it as upper-bound. Bold is best, underline is second. **For $100\rightarrow 400$ NVSR reports results of only 4 shapes in the Blender dataset. In this case, we outperform it with PSNR of $29.18$dB for our method compared to $28.5$dB of NVSR.
Figure 8: TriNeRFLet SR qualitative results. (Left) LR image; (Middle) TriNerFLet SR; (Right) Ground truth HR image.
5.3 LLFF Scene Super-Resolution
Method
@@ 5.3 LLFF Scene Super-Resolution
NeRF*
$24.47$
$0.388$
NeRF-Bi
$23.90$
$0.481$
NeRF-Liif
$24.76$
$0.292$
NeRF-Swin
$23.26$
$0.247$
NeRF-SR (SS)
25.13
$\underline{0.244}$
TriNeRFLet-SR
$\underline{25.00}$
0.203
Table 4: LLFF ×$\times 4$ Super Resolution Results for low resolution input of $378\times 504$. Bold is best, underline is second.
We turn to demonstrate the TriNeRFLet super-resolution scheme on the LLFF dataset [26], which contains 8 real-world captured scenes. As in the Blender super-resolution case, we compare our method with the same methods but for input resolution $378\times 504$ and ×$\times 4$ upscaling. We use the same settings as in the Blender $200\rightarrow 800$ experiment except for the guidance scale, which we choose to be $2.5$ instead of $7.5$. Following [40], we trained TriNeRFLet using the low-resolution version of all images and then tested the super-resolution results also on all the images. Similar to the Blender experiment, we ran LLFF TriNeRFLet super-resolution on a single A6000 RTX GPU, where each scene runs almost for $12$ hours. Table 4 reports the results. Our method is almost as good as [40] in PSNR and outperforms it in the LPIPS metric by a margin. Our LPIPS performance indicates that our method manages to create a high-resolution scene with better details and that is perceptually more similar to the high-resolution scene.

### v2

@@ 4.2 TriNeRFLet Super-Resolution (SR)
(a) TriNeRFLet based super-resolution. Low-resolution TriNeRFLet renders LR images using the LR version of the wavelet features, and fits them to the given low-resolution images. High-resolution TriNeRFLet renders HR images using all wavelet levels, and fits them to HR images from HR-set.
(b) $\text{SD}_{\text{refine}}$ process. First, a HR image is rendered. To improve its quality, noise is added to it and then it is plugged into Stable Diffusion upscaler with its LR version for conditioning. The result, which is a refined HR image is added to the HR-set.
Figure 5: TriNeRFLet Super-Resolution (SR) components.
@@
Figure 5(a) illustrates our TriNeRFLetSR strategy. First, we start by reconstructng the LR TriNeRFLet using the provided low-resolution multiview images. After training for $s_{LR}$ steps, the super-resolution process starts, and a key component of it is the stable diffusion upscaler refinement ($\text{SD}_{\text{refine}}$) step presented in Figure 5(b). Given a low-resolution ground-truth image $\bm{x}^{gt}_{LR}$, its high-resolution version $\bm{x}^{est}_{HR}$ is rendered using all wavelet levels $L$. Initially, the result will not be of high-resolution. Thus, to improve it we use a diffusion step with time $t$ randomly selected from the range $[T_{min}\;,\;T_{max}]$. Then, similar to [14], a noise with variance that depends on $t$ is added to $\bm{x}^{est}_{HR}$. Then, we plug the noisy version of $\bm{x}^{est}_{HR}$ into the SD upscaler to perform the diffusion process from step $t$ until the end, while being conditioned on the $LR$ image $\bm{x}^{gt}_{LR}$. This results in the enhanced version of $\bm{x}^{est}_{HR}$, denoted as $\bm{x}^{enhanced}_{HR}$. We denote this enhancement process by $\text{SD}_{\text{refine}}$.
We use the SD upscaler with a small modification compared to the original work [34]. In their work, the denoising step is
$\bm{x}_{t-1}=\epsilon(\bm{x}_{t},\bm{z},\varnothing)+\alpha\left(\epsilon(\bm{x}_{t},\bm{z},\bm{y})-\epsilon(\bm{x}_{t},\bm{z},\varnothing)\right),$
@@
where $\epsilon$, $\alpha$, $\bm{z}$ and $\bm{y}$ are the U-net denoiser, a guidance factor, the low-resolution image and the guidance text respectively. Instead of doing the guidance in the text direction, we do it in the low-resolution image direction and without text, i.e., $\bm{x}_{t-1}=\epsilon(\bm{x}_{t},\varnothing,\varnothing)+\alpha\left(\epsilon(\bm{x}_{t},\bm{z},\varnothing)-\epsilon(\bm{x}_{t},\varnothing,\varnothing)\right)$.
@@
Having the refined HR images, we want to use them to update the TriNeRFLet. We add these images to the set denoted HR-set. This set is initialized to be empty and gets updated sequentially. It is refreshed (i.e. emptied) every $s_{refresh}$ steps as illustrated in Algorithm 1. The images in HR-set are used as “ground-truth” images for training the HR TriNeRFLet to update all its $L$ levels.
@@
Method Mic Chair Ship Materials Lego Drums Ficus Hotdog Avg. Train Time $\downarrow$ Render FPS $\uparrow$
NeRF $32.91$ $33.00$ $28.65$ $29.62$ $32.54$ $25.01$ $30.13$ $36.18$ $31.005$ $\geq 1\text{ day}$ $\leq 0.2$
@@
Triplane $33.85$ $32.83$ $29.58$ $28.15$ $34.70$ $24.86$ $30.35$ $35.80$ $31.26$ $2$ hours $1.5$
Triplane* $34.4$ $33.5$ $29.97$ $28.1$ $35.4$ $24.9$ $30.68$ $36.16$ $31.64$ $6$ hours $1.5$
Ours: TriNeRFLet Small $35.18$ $33.76$ $30.33$ $29.14$ $35.32$ $25.66$ $33.34$ $36.24$ $32.37$ $15$ mins $1.5$
Ours: TriNeRFLet Base Light $35.77$ $35.00$ $31.10$ $29.35$ $36.44$ $25.98$ $33.96$ $36.93$ $33.07$ $1.5$ hours $1.5$
Ours: TriNeRFLet Base $\underline{36.72}$ $35.54$ $\underline{31.77}$ $29.72$ $\underline{36.66}$ $26$ $33.7$ $37.31$ $\underline{33.43}$ $6$ hours $1.5$
Ours: TriNeRFLet Large 37.22 36 32.7 30.34 37.32 26.2 $\underline{34.47}$ 37.89 34.017 $\sim 1$ day $1.5$
Table 1: Blender Dataset PSNR ($\uparrow$) results with train time and render FPS for each method. Bold is best, underline is second. For fair comparison, we also show INGP (rerun)* and Triplane*, which are trained for $6$ hours as our TriNeRFLet base.
The TriNeRFLet multiscale wavelet structure plays an important role in the SR scheme. First, its multiscale structure establishes a connection between high-resolution and low-resolution renderings by forcing them to share information (using the same first $L_{LR}$ wavelet levels). Second, wavelet regularization, especially in the layers that only the HR TriNeRFLet uses, constrains the optimization process and forces the HR TriNeRFLet to learn only useful wavelet coefficients that contribute to HR details. Thus, the combination of multiscale wavelet and diffusion models allows the method to learn the high-resolution details it exactly needs without redundancy.
The diffusion SR model that we use supports SR from $128$ to $512$. Yet, we want to support other resolutions, such as $100$ to $400$ or $200$ to $800$. For lower resolutions, e.g., going from $100$ to $400$, we pad $\bm{x}^{gt}_{LR}$ and $\bm{x}^{est}_{HR}$ with zeros in order to achieve the desired resolution before plugging them into the SD upscaler. When it finishes, the result is cropped to the original resolution. For higher resolutions, e.g., going from $200$ to $800$, we randomly crop $128$ and $512$ resolution crops from $\bm{x}^{gt}_{LR}$ and $\bm{x}^{est}_{HR}$, respectively, that correspond to the same relative location. We plug the images into the upscaler, and the result is actually an enhanced version of the crop from $\bm{x}^{est}_{HR}$. Then, we only fit the HR render to this crop instead of the whole image. When the HR-set is refreshed, a different random crop will be chosen, thus covering more areas as the process continues.
5 Experiments
We turn to present the results of our method. More visual results and ablations appear in the supplementary material. Code is available in the github repo.
5.1 3D Scene Reconstruction
[…]
Since the wavelet reconstruction component affects the runtime of TriNeRFLet, we start the training with lower resolutions of the planes, training coarse to fine. The initial plane resolution is set to $N_{base}$ and it is increased throughout the training till it reaches the size $N_{final}$ that has $L$ wavelet levels. As described above, due to the wavelet multiscale structure, when we move to the next resolution in training we simply add another level (of higher frequencies) to the wavelet representation and add it to the training, where the lower levels are kept as is (but continue to train). As mentioned in Sec. 3, the relatively high training runtime is one of TriNeRFLet framework disadvantages, and this is due to the inverse wavelet transform overhead during optimization. Thus, doing the optimization in this coarse to fine manner mitigates this drawback and accelerates the training (inverse wavelet on smaller planes is faster).
To evaluate the performance of our method, we use the Blender dataset [28], which contains $8$ synthetic 3D scenes with complex geometry and lighting. We compare TriNeRFLet with the vanilla NeRF [28], regular Triplane (with base config), INGP [29] and 3D Gaussian Splatting [21]. Results are reported in Table 1. For INGP we report 3 versions: INGP (paper) denotes the officially published results by [29], and INGP (rerun) and INGP (rerun)* denote the results that we obtained by running INGP (using [29] software) for $5$ mins and $6$ hours respectively. Also, Triplane and Triplane* denote Triplane results when trained for $5$ mins and $6$ hours respectively. We ran all TriNeRFLet versions on a single A6000 GPU. The reported results indeed demonstrate the improvement in performance that the multiscale wavelet brings, as it outperforms regular Triplane by a significant margin, and improves over the current SOTA, even in case they were trained for a longer time.
Regarding rendering time, the TriNeRFLet (all versions) frames per second (FPS) are approximately $1.5$, while INGP and 3D Gaussian Splatting are approximately $1$ and $135$ respectively. These FPS are tested on a single A6000 GPU. Note that TriNeRFLet rendering FPS is compatible and even slightly better than INGP, which is known to be fast within the NeRF approaches. 3D Gaussian Splatting, which is not a NeRF method, is faster in its rendering time since it uses a different internal structure than NeRF. When compared only to NeRF-based solutions, our method is very fast with competitive reconstruction performance. While our rendering time is faster, the training time of TriNeRFLet is higher than INGP. This is due to the wavelet reconstruction, which needs to be done at each training step but only once during rendering time.
5.2 3D Scene Super-Resolution
Blender. We turn to present our results for TriNeRFLet SR. The objective here is to generate novel views with ×$\times 4$ resolution than the input images. To this end, we tested two different settings on the Blender dataset [28], $100$ to $400$ and $200$ to $800$. For the first one, we use $N_{LR}=256$ and $N=1024$, while in the second setting $N_{LR}=512$ and $N=2048$. Further technical details appear in the Supplementary Materials.
We compare our method with NeRF*, NeRF-Bi, NeRF-Liif, NeRF-Swin, NeRF-SR (SS) (these methods do not require 3D supervision and are described in [41]) and NVSR (results of $200\rightarrow 800$ were provided to us by the authors) [4]. NeRF* is regular NeRF trained on LR views and then rendered in HR resolution, NeRF-Bi, NeRF-Liif and NeRF-Swin are trained on LR views, then rendered in LR resolution and upscaled using different 2D SR methods [41].
Results are reported in Table 3. Note that NVSR requires 3D supervision (paired LR and HR multiview images) and therefore it is considered only as a reference. When compared only to methods that use 2D supervised methods, we perform the best in most of the cases. Also, our method’s results are very close to NVSR, and even slightly outperforms it in some cases, which demonstrates the superiority of our method despite the fact that it was not trained using 3D data. Qualitative examples of our approach appear in Figure 6 and sup. mat.
Impact of multiscale on SR. To demonstrate the impact of TriNeRFLet multiscale on the SR scheme, we modify it to neutralize the multiscale element. This is done by replacing the low-resolution rendering that uses low-resolution wavelet features (first $L_{LR}$ layers) with a downscaled HR rendering that uses all wavelet layers (i.e. without multiscale render). In the same manner, INGP is also tested as a representation. Results in Table 2 demonstrate the important contribution of the TriNeRFLet multiscale structure to the SR scheme, utilizing its full potential, and that without it the SR scheme performs worse.
Impact of LPIPS loss. To demonstrate the LPIPS loss impact on SR, we trained the scheme without the LPIPS loss. Results are reported in Table 2, and as noticed, the LPIPS component indeed improves the SR scheme performance.
TriNeRFLetSR TriNeRFLetSR w/o multiscale TriNeRFLetSR w/o LPIPS Our SR scheme with INGP
@@ 5 Experiments
[…]
Table 2: Ablation: Impact of wavelet multiscale and LPIPS on SR. To show the significance of having the multi-scale structure of the wavelet, we apply our SR scheme without using the multiscale rendering or with INGP instead of TriNeRFLet rendering. We also show the significance of adding the LPIPS loss.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
       LR image           NVSR    TriNeRFLet SR   HR Ground Truth
Figure 6: TriNeRFLet SR qualitative results. Note the visual improvement that is achieved by our approach to NeRF SR compared to LR reconstruction. Moreover, the qualitative results of our method compete with NVSR, which unlike our strategy is a 3D supervised method that we consider as a reference. In the sup. mat. we provide addtional visual examples with more comparisons to other methods.
Blender Blender LLFF
$100\rightarrow 400$ $200\rightarrow 800$ $378\times 504\rightarrow 1512\times 2016$
[…]
NeRF* $25.56$ $0.170$ $0.881$ $27.47$ $0.128$ $0.900$ $24.47$ $0.388$ $0.701$
NeRF-Bi $24.74$ $0.244$ $0.868$ $26.67$ $0.175$ $0.900$ $23.90$ $0.481$ $0.676$
NeRF-Liif $25.36$ $0.125$ $0.885$ $27.34$ $0.096$ $0.912$ $24.76$ $0.292$ $0.723$
NeRF-Swin $24.85$ $0.108$ $0.881$ $26.78$ $0.086$ $0.906$ $23.26$ $0.247$ $0.685$
NeRF-SR (SS) $\underline{28.07}$ $\underline{0.071}$ 0.921 $28.46$ $0.076$ $0.921$ 25.13 $\underline{0.244}$ $\underline{0.730}$
TriNeRFLetSR 28.55 0.061 $\underline{0.913}$ $\underline{29.49}$ 0.051 $\underline{0.930}$ $\underline{25.00}$ 0.203 0.771
NVSR (3D supervised) **∗ ∗ 29.53 $\underline{0.054}$ 0.931 **∗ ∗
Table 3: Super Resolution ×$\times 4$ Results. NVSR is trained with multiview (3D) supervision and therefore we consider it as reference. Bold is best, underline is second. **For $100\rightarrow 400$ NVSR reports results of only 4 shapes in the Blender dataset. In this case, we outperform it with PSNR of $29.18$dB for our method compared to $28.5$dB of NVSR. For LLFF, NVSR did not report any results on this resolution or LLFF average.
LLFF. We turn to demonstrate the TriNeRFLet-based SR scheme on the LLFF dataset [27], which contains 8 real-world captured scenes. As in the Blender SR case, we compare our method with the same methods but for input resolution $378\times 504$ and ×$\times 4$ upscaling. Following [41], we trained TriNeRFLet using the low-resolution version of all images and then tested the SR results also on all the images. Table 3 reports the results. Our method is almost as good as [41] in PSNR and outperforms it in the LPIPS and SSIM metrics by a margin. Our LPIPS and SSIM performance indicates that our method manages to create a scene with better visual details and that is perceptually more similar to the ground-truth high-resolution scene.